OpenStudy (anonymous):

Consider the reaction below. If you start with 2.00 moles of C3H8 (propane) and 2.00 moles of O2, how many moles of carbon dioxide can be produced? C3H8(g) + 5O2(g) 3CO2(g) + 4H2O(g) 8.00 1.20 6.00 2.00

OpenStudy (anonymous):

so there is a 1:5 ratio between the reactants, whereas you are provided 2 moles of each. So what is the limiting reactant? If you said oxygen you would be correct! 2 moles of O2 are used and 0.4 moles of C3H8 are used (2/5= .4). Now we can look at the moles ratio of oxygen to carbon dioxide: 5:3. Meaning for every 5 moles of oxygen reacting, we yield 3 moles of carbon dioxide. So if 2 moles react: 2*(3/5)= 1.2 moles Let me know if this is clear!

OpenStudy (anonymous):

Yes This is very clear! thank you. sorry for my late response IO was completing other problems! But how would you go about solving the rest of this problem?

OpenStudy (anonymous):

Hi Josephine, So step by step you go from: (1) given moles and mole ratio of reactant, find the limiting reactant (2) the limiting reactant will FULLY react, meaning there will be no moles of it left after the rxn has gone to completion (3) Since oxygen is the limiting reactant, 2 moles of oxygen will be used up (4) now look at the mole ratio of the oxygen to carbon dioxide (5:2), this means for every 5 moles of oxygen, you will produce 2 moles of carbon dioxide (5) so the math goes (2 mole O2) x (3 mole CO2)/(5 mole O2) = 6/5 = 1.2 mole CO2 if there is more to the problem let me know!